An economist reports that 238 out of a sample of 700 middle-income American households actively participate in the stock market. [You may find it useful to reference the z table.]
  

a. Construct the 95% confidence interval for the proportion of middle-income Americans who actively participate in the stock market. (Round intermediate calculations to at least 4 decimal places. Round "z" value and final answers to 3 decimal places.)  



b. Can we conclude that the percentage of middle-income Americans who actively participate in the stock market is not 38%?


  • Yes, since the confidence interval contains the value 0.38.

  • Yes, since the confidence interval does not contain the value 0.38.

  • No, since the confidence interval contains the value 0.38.

  • No, since the confidence interval does not contain the value 0.38.
Answer #1

solution? Given that, n- 700 X 238 P = x = 238 - 0.340 700 point estimate =0.340 L-=0.340 = 0.660 2 At as o% confidence levelNo since the confidence interval does not contain the value 0.38 option CD) is correct